OpenStudy (anonymous):

How do i know if a metod of choice problem is many solutions, no solutions, or none for -7x+7y=7 and 2x-2y=-18??

OpenStudy (anonymous):

7y = 7x + 7 y = x + 1 2x - 2y = 18 2y = 2x - 18 y = x - 9 they are parallel and with different y-intercepts. Therfore, it's no solution/intersection